write an equivalent equation using polar coordinates: x^2=6y
Answer #1

To convert rectangular coordinates into polar coordinates put x=rcos(theta) y=rsin(theta)

Given x^2=6y

(rcos(theta))^2= 6rsin(theta)

r^2cos^2(theta)= 6rsin(theta)

cancelling r on both sides

rcos^2(theta)=6sin(theta)

r=(6sin(theta))/(cos^2(theta))

r=6(sin(theta)/cos(theta))(1/cos(theta))

r=6tan(theta)sec(theta)